# A Study of Fairness Concerns in AI-based Mobile App Reviews

### v2

@@ Abstract.
Fairness is one of the socio-technical concerns that must be addressed in AI-based systems. Unfair AI-based systems, particularly unfair AI-based mobile apps, can pose difficulties for a significant proportion of the global population. This paper aims to analyze fairness concerns in AI-based app reviews. We first manually constructed a ground-truth dataset, including a statistical sample of fairness and non-fairness reviews. Leveraging the ground-truth dataset, we developed and evaluated a set of machine learning and deep learning classifiers that distinguish fairness reviews from non-fairness reviews. Our experiments show that our best-performing classifier can detect fairness reviews with a precision of 94%. We then applied the best-performing classifier on approximately 9.5M reviews collected from 108 AI-based apps and identified around 92K fairness reviews. Next, applying the K-means clustering technique to the 92K fairness reviews, followed by manual analysis, led to the identification of six distinct types of fairness concerns (e.g., ‘receiving different quality of features and services in different platforms and devices’ and ‘lack of transparency and fairness in dealing with user-generated content’). Finally, the manual analysis of 2,248 app owners’ responses to the fairness reviews identified six root causes (e.g., ‘copyright issues’) that app owners report to justify fairness concerns.
@@ 1. Introduction
The global AI market size, particularly in the form of Machine Learning (ML) and Deep Learning (DL), is predicted to exceed USD 1.6 trillion by 2030 (mar, 2022). The applications of AI technologies in our modern life and society vary, ranging from detecting diseases to predicting traffic jams, from enabling personalized shopping to identifying important targets in combat environments (Mehrabi et al., 2021). Despite the wide adoption of AI solutions, a growing concern exists about ensuring the responsible behavior of AI solutions (Mikalef et al., 2022). The increasing need for responsible AI systems arises from the possibility that they could produce biased results, have errors, or lack sufficient transparency (Trocin et al., 2021). Hence, organizations and governments worldwide have started developing principles and guidelines on the responsible use and adoption of AI (Dignum, 2022). Fairness is a key element of responsible AI (Ahmad et al., 2023). There is no single definition of fairness in AI systems, and varies in different domains (Ekstrand et al., 2022). In AI-based algorithmic decision-making, fairness is defined as “the absence of any prejudice or favoritism towards an individual or group based on their inherent or acquired characteristics” (Mehrabi et al., 2021). Others have a broader view and see achieving fairness in AI systems as a socio-technical challenge (Ekstrand et al., 2022; Selbst et al., 2019; Bird et al., 2020; Lu et al., 2022). This means that fairness concerns in AI systems cannot only stem from technical components and data but can also be attributed to the process of building AI systems, humans, and governance (e.g., decisions made by developers or policies adopted by providers). However, a growing number of software systems leveraging AI solutions (AI-based software systems) have been accused of being unfair and producing biased and discriminatory outcomes, causing some (major) difficulties for individuals, society, and businesses (Mehrabi et al., 2021; Galhotra et al., 2017).
@@
The growth and popularity of AI have also triggered a wave of mobile software applications utilizing AI technologies (e.g., ML/DL technologies). Such a wave has been further energized by the emergence of mobile AI frameworks and libraries such as TensorFlow Lite that enable executing AI tasks solely on smartphones (Xu et al., 2019). We refer to mobile software applications that utilize AI technologies as AI-based mobile applications (apps). With 6.64 billion smartphone users worldwide (83.32% of the current world population) (Turner, 2023), mobile software apps are the most prevailing type of software system used by a wide range of individuals and groups with different characteristics (e.g., age, education level, cultural background, race, and gender) (Sarker, 2019; Fazzini et al., 2022; Khalajzadeh et al., 2022). While unfair behaviours and outcomes can manifest in any mobile application, including those not powered by AI, we posit that the chance of unfair behaviours and outcomes in AI-based mobile apps may be higher than in non-AI-based apps (Mehrabi et al., 2021; Mikalef et al., 2022; Madaio et al., 2022; Lewicki et al., 2023). This increased risk is attributed to the complex nature of AI technologies, which can inadvertently encode and perpetuate biases present in their training data or opaque decision-making algorithms (Ahmad et al., 2023; Nguyen-Duc et al., 2023). Moreover, considering the pervasiveness of mobile apps, unfair behaviours and outcomes in (AI-based) mobile apps have far-reaching implications, affecting a significant proportion of the global population. Lastly, to the best of our knowledge, no work has focused on providing an accurate and fine-grained view of different types of fairness concerns and their possible root causes in AI-based apps. Having said that, our work in this paper sets the stage for future research, advocating for a broader application of our dataset curation approach and classifier (discussed later in the paper) to explore and mitigate unfair behaviors in the wider landscape of mobile applications, AI-based or otherwise.
@@
This research aims to investigate fairness concerns in AI-based app reviews. To this end, we collected $\approx$9.5M reviews from 108 Android AI-based apps from the Google Play Store. We then manually constructed a dataset (ground truth) composed of fairness and non-fairness reviews. In this study, a fairness review is a review that discusses a fairness concern (topic). Fig. 1 shows an example of a fairness review and the response made by the app owner to justify the raised fairness concern. Next, we evaluated a set of ML and DL binary classifiers against the ground-truth dataset, which aims to distinguish fairness reviews from non-fairness reviews. Our experiments show that our best-performing classifier can detect fairness reviews with a precision of 94%. Although the binary classifiers are promising in identifying fairness reviews, they do not provide a fine-grained view of different fairness concerns. Hence, we developed and experimented with the K-means clustering technique, followed by a manual analysis to cluster and summarize fairness reviews. Our study shows that users discuss the top six fairness concerns in AI-based app reviews: ‘receiving different quality of features and services in different platforms and devices’, ‘feeling linguistic discrimination, ‘lack of transparency and fairness in dealing with user-generated content’, ‘feeling gender and racial discrimination’, ‘feeling biased censorship and promotion’, and ‘unfair and non-transparent advertisement and subscription policies’. The app owners can use our approach in practice to automatically delineate key fairness-related concerns (and any underlying reasons) in their apps, combining our binary classifier and the clustering approach. To understand the underlying reasons for fairness concerns, we manually analyzed 2,248 responses made by app owners to the fairness reviews. This led to the identification of six root causes of the fairness concerns: ‘copyright issues’, ‘development complexity’, ‘buggy code’, ‘external factors’, ‘development cost’, and ‘user usage and awareness’. The primary contributions of this paper are:
@@
  • We develop ML and DL binary classifiers that accurately distinguish fairness from non-fairness reviews.
@@
Paper Organization. Section 2 summarizes the related literature. Section 3 introduces our research questions and describes our dataset curation process with fairness and non-fairness reviews. Section 4 elaborates on the experiments with a set of binary classifiers that detect fairness concerns in app reviews, followed by a classification of fairness reviews using the best-performing classifier and the K-means clustering technique to summarize fairness concerns for answering RQ1 in Section 5. We report the findings of RQ2 about the root causes of fairness concerns in Section 6. Section 7 discusses the main findings. Section 8 outlines the threats to validity, and Section 9 concludes this work.
@@ 2.2. Human, Social, and Ethical Concerns in App Reviews
Human and ethical values in mobile app development have been the focus of some studies (e.g., (Obie et al., 2022; Shams et al., 2020)). Obie et al. (Obie et al., 2022) focused on ‘honesty’ as one of the human values and identified 10 types of honesty violations (e.g., ‘delusive subscriptions’, ‘false advertainments’) discussed by users in mobile app reviews. They further developed several ML classifiers to automatically detect honesty violations in app reviews. Shahin et al. (Shahin et al., 2023) studied app reviews to understand the gender-related requirements and concerns discussed by users. To this end, they built a binary classifier to identify reviews that include a gender discussion, followed by a qualitative study. They argued that such gender-related requirements and concerns are essential to building gender-inclusive apps. Khalajzadeh et al. (Khalajzadeh et al., 2022) constructed a taxonomy that categorizes human-centric issues that users face when working with mobile apps into three high-level groups: “App Usage”, “Inclusiveness”, and “User Reaction”. They then developed several ML/DL classifiers that automatically detect and categorize these three human-centric issues from app reviews. Note that the studies summarized above have either focused on fairness from a technical perspective (learning algorithms and data) or investigated socio-technical aspects such as privacy and honesty other than fairness in app reviews.
@@
Fairness concerns have been only recently investigated in app reviews by Arony et al. (Arony et al., 2023) as part of the construction of a taxonomy for inclusiveness. The taxonomy was developed based on inclusiveness-related discussions in Twitter, Reddit, and Google Play Store, which includes six major categories: “Fairness”, “Technology”, “Privacy”, “Demography”, “Usability”, and “Other Human Values”. The “Fairness” category has three sub-categories “Terms/Conditions”, “Recommendation”, and “Services”. Arony et al. also leveraged five DL classifiers to automatically distinguish inclusiveness-related discussions from non-inclusiveness-related discussions. These classifiers do not dedicatedly aim to detect fairness reviews. We argue that there is still a significant gap in understanding the types of fairness concerns raised by users in AI-based mobile app reviews and their potential root causes from the app owners’ perspective. In this research, we aim to address this gap with an exhaustive analysis of 108 AI-based apps, specifically analyzing user reviews to extract fairness concerns. Our research not only adds a new dimension to this existing body of research but also brings us a step closer to comprehending fairness from the users’ viewpoint.
@@ 3. Research Questions and Dataset
This paper aims to deeply analyze the fairness of AI-based apps. This will be achieved by answering the following overarching research questions (RQs).
RQ1: What types of fairness concerns are raised by users in AI-based mobile app reviews?
Motivation. The users share a wide range of information and opinions on app stores, including socio-technical issues they experience when using mobile apps (Wang et al., 2022). The richness of feedback shared by users has motivated researchers and app providers to mine app reviews, leading to extracting actionable information for app owners and developers. Currently, there is a lack of understanding of the types of fairness concerns that users may raise in AI-based mobile apps. We argue that user reviews of AI-based mobile apps can be a rich resource to uncover fairness concerns from the users’ perspective. However, manually identifying and classifying such fairness concerns in millions of app reviews is challenging. While automated techniques, including ML and DL classifiers, have been proposed in recent years for mining app reviews (e.g., (Wang et al., 2022; Nema et al., 2022; Obie et al., 2022; Khalajzadeh et al., 2022)), none of these automated techniques has focused on fairness concerns. Furthermore, applying general-purpose automated techniques may misidentify and mis-classify fairness concerns. To help close this gap, we plan to experiment with several ML and DL binary classifiers and a clustering technique that detect and summarize fairness concerns from the user reviews of AI-based apps.
RQ2: What are the root causes of fairness concerns in apps, based on the justification by app owners?
@@ 3.1. Dataset Curation
There is no current dataset (to the best of our knowledge) with fairness and non-fairness app reviews. We curated a dataset to evaluate our binary classification and clustering techniques. Fig. 2 illustrates our dataset curation process, discussed next.
@@ 3.1.1. Identification of AI-based apps
Several studies have developed approaches to identify AI-based apps in Google Play Store. To have a relatively comprehensive list of AI-based apps, we used the existing datasets of AI-based apps (Xu et al., 2019; Li et al., 2022a). Xu et al. (Xu et al., 2019) identified 211 offline-mode DL Android apps in Google Play Store by examining app installation packages. As DL is only one sub-category of AI techniques, Li et al. (Li et al., 2022a) extended this to develop an approach called AI Discriminator that works based on app dissection and keyword matching and identifies apps that use any AI frameworks. Li et al. collected 56,682 AI-based apps in the AndroZoo repository (i.e., a collection of Android apps (Allix et al., 2016)). The two datasets had 56,893 AI-based apps in total. We removed duplicates from these two datasets and removed apps with less than 10K reviews, i.e., only included ‘popular’ apps with sufficient user feedback. This process led to 108 AI-based apps. These apps come from 23 distinct categories (e.g., social, communication, and productivity). Our replication package (Nasab et al., 2024) provides a demographic of these apps.

### v3

# Fairness Concerns in App Reviews: A Study on AI-based Mobile Apps

@@ Abstract.
Fairness is one of the socio-technical concerns that must be addressed in AI-based systems. Unfair AI-based systems, particularly unfair AI-based mobile apps, can pose difficulties for a significant proportion of the global population. This paper aims to analyze fairness concerns in AI-based app reviews. We first manually constructed a ground-truth dataset, including 1,132 fairness and 1,473 non-fairness reviews. Leveraging the ground-truth dataset, we developed and evaluated a set of machine learning and deep learning models that distinguish fairness reviews from non-fairness reviews. Our experiments show that our best-performing model can detect fairness reviews with a precision of 94%. We then applied the best-performing model on approximately 9.5M reviews collected from 108 AI-based apps and identified around 92K fairness reviews. Next, applying the K-means clustering technique to the 92K fairness reviews, followed by manual analysis, led to the identification of six distinct types of fairness concerns (e.g., ‘receiving different quality of features and services in different platforms and devices’ and ‘lack of transparency and fairness in dealing with user-generated content’). Finally, the manual analysis of 2,248 app owners’ responses to the fairness reviews identified six root causes (e.g., ‘copyright issues’) that app owners report to justify fairness concerns.
@@ 1. Introduction
The global AI market size, particularly in the form of Machine Learning (ML) and Deep Learning (DL), is predicted to exceed USD 1.6 trillion by 2030 (mar, 2022). The applications of AI technologies in our modern life and society vary, ranging from detecting diseases to predicting traffic jams, from enabling personalized shopping to identifying important targets in combat environments (Mehrabi et al., 2021). Despite the wide adoption of AI solutions, a growing concern exists about ensuring the responsible behavior of AI solutions (Mikalef et al., 2022). The increasing need for responsible AI systems arises from the possibility that they could produce biased results, have errors, or lack sufficient transparency (Trocin et al., 2021). Hence, organizations and governments worldwide have started developing standards, frameworks, and guidelines on the responsible use and adoption of AI (Dignum, 2022; Team, 2024; Sameki, 2024; IEEE, 2021). Fairness is one of the important principles in the existing ethical and responsible AI frameworks and standards. For example, Microsoft’s Responsible AI framework includes fairness as one of its six core principles, along with reliability and safety, privacy and security, inclusiveness, transparency, and accountability (Sameki, 2024). Similarly, Google emphasizes fairness in its responsible AI practices to ensure that AI systems are developed to enhance people’s lives in a fair manner (Team, 2024). There is no single definition of fairness in AI systems, and varies in different domains (Ekstrand et al., 2022). In AI-based algorithmic decision-making, fairness is defined as “the absence of any prejudice or favoritism towards an individual or group based on their inherent or acquired characteristics” (Mehrabi et al., 2021). Others have a broader view and see achieving fairness in AI systems as a socio-technical challenge (Ekstrand et al., 2022; Selbst et al., 2019; Bird et al., 2020; Lu et al., 2022). This means that fairness concerns in AI systems cannot only stem from technical components and data but can also be attributed to the process of building AI systems, humans, and governance (e.g., decisions made by developers or policies adopted by providers). However, a growing number of software systems leveraging AI solutions (AI-based software systems) have been accused of being unfair and producing biased and discriminatory outcomes, causing some (major) difficulties for individuals, society, and businesses (Mehrabi et al., 2021; Galhotra et al., 2017).
@@
The growth and popularity of AI have also triggered a wave of mobile software applications utilizing AI technologies (e.g., ML/DL technologies). Such a wave has been further energized by the emergence of mobile AI frameworks and libraries such as TensorFlow Lite that enable executing AI tasks solely on smartphones (Xu et al., 2019). Inspired by (Li et al., 2022a; Xu et al., 2019), we define AI-based mobile applications (apps) as mobile apps that utilize AI frameworks, including ML-based, DL-based, and AI service-based frameworks (e.g., TensorFlow Lite (Ten, 2024), Google AI (Goo, 2024)). With 6.64 billion smartphone users worldwide (83.32% of the current world population) (Turner, 2023), mobile software apps are the most prevailing type of software system used by a wide range of individuals and groups with different characteristics (e.g., age, education level, cultural background, race, and gender) (Sarker, 2019; Fazzini et al., 2022; Khalajzadeh et al., 2022). While unfair behaviours and outcomes can manifest in any mobile app, including those not powered by AI, we posit that the chance of unfair behaviours and outcomes in AI-based mobile apps may be higher than in non-AI-based apps (Mehrabi et al., 2021; Mikalef et al., 2022; Lewicki et al., 2023). This increased risk is attributed to the complex nature of AI technologies, which can inadvertently encode and perpetuate biases present in their training data or opaque decision-making algorithms (Mehrabi et al., 2021; Belenguer, 2022; Lepri et al., 2018). Moreover, considering the pervasiveness of mobile apps, unfair behaviours and outcomes in (AI-based) mobile apps have far-reaching implications, affecting a significant proportion of the global population. Lastly, to the best of our knowledge, no work has focused on providing an accurate and fine-grained view of different types of fairness concerns and their possible root causes discussed in AI-based app reviews. Having said that, our work in this paper sets the stage for future research, advocating for a broader application of our dataset curation approach and classifier (discussed later in the paper) to explore and mitigate unfair behaviors in the wider landscape of mobile apps, AI-based or otherwise.
@@
This research aims to investigate fairness concerns raised in mobile app reviews. We selected AI-based mobile apps as the subjects of our study for the reasons outlined earlier. To this end, we first collected $\approx$9.5M reviews from 108 Android AI-based apps from the Google Play Store. We then manually constructed a dataset (ground truth) composed of 1,132 fairness and 1,473 non-fairness reviews. We define a fairness review as a review in which users express that they are treated differently, whether intentionally or unintentionally, due to their inherent, acquired, or context-specific attributes (e.g., age, language, location, gender) by AI-based apps or the decisions and policies of the AI-based app owners. Additionally, a fairness review can reflect users’ perceptions that the outcomes of AI-based apps, or the decisions and policies of the app owners, are inconsistent, non-transparent, and unreasonable. Fig. 1 shows an example of a fairness review and the response made by the app owner to justify the raised fairness concern. Next, we evaluated a set of ML and DL models against the ground-truth dataset, which aims to distinguish fairness reviews from non-fairness reviews. Our experiments show that our best-performing model can detect fairness reviews with a precision of 94%. Although the models are promising in identifying fairness reviews, they do not provide a fine-grained view of different fairness concerns. Hence, we developed and experimented with the K-means clustering technique, followed by a manual analysis to cluster and summarize fairness reviews. Our study shows that users discuss the top six fairness concerns in AI-based app reviews: ‘receiving different quality of features and services in different platforms and devices’, ‘feeling linguistic discrimination, ‘lack of transparency and fairness in dealing with user-generated content’, ‘feeling gender and racial discrimination’, ‘feeling biased censorship and promotion’, and ‘unfair and non-transparent advertisement and subscription policies’. The app owners can use our approach in practice to automatically delineate key fairness-related concerns (and any underlying reasons) in their apps, combining our model and the clustering approach. To understand the underlying reasons for fairness concerns, we manually analyzed 2,248 responses made by app owners to the fairness reviews. This led to the identification of six root causes of the fairness concerns: ‘copyright issues’, ‘development complexity’, ‘buggy code’, ‘external factors’, ‘development cost’, and ‘user usage and awareness’. The primary contributions of this paper are:
@@
  • We manually construct a ground-truth dataset consisting of 1,132 fairness reviews and 1,473 non-fairness reviews.
  • We develop ML and DL models that accurately distinguish fairness from non-fairness reviews.
@@
Paper Organization. Section 2 summarizes the related literature. Section 3 introduces our research questions and describes our dataset curation process with fairness and non-fairness reviews. Section 4 elaborates on the experiments with a set of models that detect fairness reviews in app reviews. Section 5 applies the K-means clustering technique on fairness reviews to identify fairness concerns. We report the root causes of fairness concerns in Section 6. Section 7 discusses the main findings. Section 8 outlines the threats to validity, and Section 9 concludes this work.
@@ 2.1. Fairness in AI Systems
Several literature reviews and surveys have been conducted on fairness in AI. Fabris et al. (Fabris et al., 2022) studied over 200 datasets used in fairness research in different domains (e.g., health, natural sciences) to develop standardized and searchable documentation for algorithmic fairness datasets. This effort aims to alleviate the documentation debt present in existing fairness datasets. Hort et al. (Hort et al., 2023) conducted a survey of bias mitigation techniques for ML classifiers. They found that most of the bias mitigation techniques proposed in the literature attempt to mitigate bias during training ML models (i.e., in-processing mitigation techniques), followed by pre-processing mitigation techniques (i.e., those that reduce bias by changing the training dataset). Post-processing is another type of bias mitigation technique that is less explored in the literature. Such techniques are applied after a classifier has been successfully trained. In a review study, Mehrabi et al. (Mehrabi et al., 2021) found that bias and unfairness primarily originate from data and algorithms. Their study also developed a taxonomy of fairness definitions. Another survey by Caton and Haas (Caton and Haas, 2024) revealed that most current attempts focus on improving fairness in supervised binary classification. Additionally, they summarized the methods proposed in the literature to measure fairness in ML.
@@ 2.2. Human, Social, and Ethical Concerns in App Reviews
Human and ethical values in mobile app development have been the focus of some studies (e.g., (Obie et al., 2022; Shams et al., 2020)). Shahin et al. (Shahin et al., 2023) studied app reviews to understand the gender-related requirements and concerns discussed by users. To this end, they built a binary classifier to identify reviews that include a gender discussion, followed by a qualitative study. They argued that such gender-related requirements and concerns are essential to building gender-inclusive apps. Khalajzadeh et al. (Khalajzadeh et al., 2022) constructed a taxonomy that categorizes human-centric issues that users face when working with mobile apps into three high-level groups: “App Usage”, “Inclusiveness”, and “User Reaction”. They then developed several ML/DL classifiers that automatically detect and categorize these three human-centric issues from app reviews. Note that the studies summarized above have either focused on fairness from a technical perspective (learning algorithms and data) or investigated socio-technical aspects such as privacy and honesty other than fairness in app reviews.
@@
Fairness concerns have been only recently investigated in app reviews by Arony et al. (Arony et al., 2023) as part of the construction of a taxonomy for inclusiveness. The taxonomy was developed based on inclusiveness-related discussions in Twitter, Reddit, and Google Play Store, which includes six major categories: “Fairness”, “Technology”, “Privacy”, “Demography”, “Usability”, and “Other Human Values”. The “Fairness” category has three sub-categories “Terms/Conditions”, “Recommendation”, and “Services”. Arony et al. also leveraged five DL classifiers to automatically distinguish inclusiveness-related discussions from non-inclusiveness-related discussions. Obie et al. (Obie et al., 2022) focused on ‘honesty’ as one of the human values and manually identified 10 types of honesty violations (e.g., ‘delusive subscriptions’, ‘false advertainments’) discussed by users in mobile app reviews. They further developed several ML classifiers to automatically classify user reviews into honesty violation reviews and non-honesty violation reviews. Two types of the identified honesty violations are related to fairness: ‘unfair cancellation and refund policies’ and ‘unfair fees’, as the definition of honesty they considered was broad and encompassed fairness, too. While Obie et al. (Obie et al., 2022) and Arony et al. (Arony et al., 2023) had touched upon some aspects of fairness within their respective scopes of honesty and inclusiveness, their research did not focus on creating a dedicated approach for identifying fairness reviews, automatically classifying types of fairness concerns, or pinpointing the root causes of these concerns. We argue that there is still a significant gap in understanding the types of fairness concerns raised by users in AI-based mobile app reviews and their potential root causes from the app owners’ perspective. In this research, we aim to address this gap with an exhaustive analysis of 108 AI-based apps, specifically analyzing user reviews to extract fairness concerns. Our research not only adds a new dimension to this existing body of research but also brings us a step closer to comprehending fairness from the users’ viewpoint.
@@ 3. RESEARCH QUESTIONS AND DATASET
This paper aims to deeply analyze fairness in AI-based app reviews. This will be achieved by answering the following overarching research questions (RQs).
RQ1: Can we effectively identify fairness reviews in AI-based app reviews?
Motivation. The users share a wide range of information and opinions on app stores, including socio-technical issues they experience when using mobile apps (Wang et al., 2022). The richness of feedback shared by users has motivated researchers and app providers to mine app reviews, leading to extracting actionable information for app owners and developers. We argue that user reviews of AI-based mobile apps can be a rich resource for uncovering fairness concerns from the users’ perspective. However, manually identifying fairness reviews in millions of app reviews is challenging. While automated techniques, including ML and DL models, have been proposed in recent years for mining app reviews (e.g., (Wang et al., 2022; Nema et al., 2022; Obie et al., 2022; Khalajzadeh et al., 2022)), none of these automated techniques has focused on fairness concerns. Furthermore, applying general-purpose automated techniques may misidentify and misclassify fairness concerns. To help close this gap, we plan to experiment with several models to automatically distinguish fairness reviews from non-fairness reviews.
RQ2: What types of fairness concerns are raised by users in AI-based app reviews?
Motivation. Distinguishing fairness reviews from non-fairness reviews is the first step to understanding fairness reviews. Still, it cannot provide a fine-grained view of different types of fairness concerns raised by various users of AI-based apps. Hence, this RQ aims to use a clustering technique to summarize the fairness concerns raised in the fairness reviews identified from RQ1.
RQ3: What are the root causes of fairness concerns in apps, based on the justification by app owners?
@@ 3.1. Dataset Curation
There is no current dataset (to the best of our knowledge) with fairness and non-fairness app reviews. We curated a dataset to evaluate our learning models and clustering techniques. Fig. 2 illustrates our dataset curation process, discussed next.
@@
Table 1. The number of the identified AI-based apps in each category.
# Categories Number of AI Apps # Categories Number of AI Apps
1 Photography 15 13 Health & Fitness 3
2 Productivity 13 14 Food & Drink 3
3 Communication 10 15 Lifestyle 3
4 Shopping 10 16 Role Playing 2
5 Tools 9 17 Arcade 1
6 Social 7 18 Education 1
7 Video Players & Editors 6 19 Strategy 1
8 Travel & Local 5 20 Art & Design 1
9 Entertainment 4 21 Auto & Vehicles 1
10 Finance 4 22 Puzzle 1
11 Business 4 23 Personalization 1
12 Music & Audio 3
@@ 3.1.1. Identification of AI-based apps
Several studies have developed approaches to identify AI-based apps in Google Play Store. To have a relatively comprehensive list of AI-based apps, we used the existing datasets of AI-based apps (Xu et al., 2019; Li et al., 2022a). Xu et al. (Xu et al., 2019) identified 211 offline-mode DL Android apps in Google Play Store by examining app installation packages. They defined DL Android apps as mobile apps using one of the 16 popular DL frameworks such as TensorFlow Lite (Ten, 2024), Caffe2 (caf, 2024), etc. As DL is only one sub-category of AI techniques, Li et al. (Li et al., 2022a) extended this to develop an approach called AI Discriminator that works based on app dissection and keyword matching and identifies apps that use any AI frameworks. Li et al. collected 56,682 AI-based apps in the AndroZoo repository (i.e., a collection of Android apps (Allix et al., 2016)). They defined AI-based apps as apps that use any of 33 AI frameworks categorized into classic ML-based, DL-based, and AI service-based frameworks (e.g., Google AI (Goo, 2024)). The two datasets had 56,893 AI-based apps in total. We removed duplicates from these two datasets and removed apps with less than 10K reviews, i.e., only included ‘popular’ apps with sufficient user feedback. This process led to 108 AI-based apps. These apps come from 23 distinct categories (e.g., social, communication, and productivity). Table 1 shows the number of AI-based apps in each category. Our replication package (Nasab et al., 2024) provides a demographic of these apps.